OpenStudy (anonymous):

Find the general solution to (b) dy/dt = t^2y^3

OpenStudy (amistre64):

this thing slow, or is it just me?

OpenStudy (anonymous):

\[\frac{dy}{y^3}=t^2dt\]Integrate both sides to get:\[\frac{-1}{2y^2}=\frac{t^3}{3}+C\]Solve for y in terms of t.

OpenStudy (amistre64):

separables: \[\int \left(y^{-3}dy=t^2dt\right)\]

OpenStudy (anonymous):

what does y =?

OpenStudy (amistre64):

thats what solving it will actually get you. If you cant remember how to integrate simple powers, then you might need to go back and review that.